# Recov-Vision: Linking Street View Imagery and Vision-Language Models for Post-Disaster Recovery

### v1

@@ Acknowledgments
Appendix A VLM Prompts and Few-shot Examples
@@
house_destruction”: true/false,
structural_damage”: true/false,
exterior_debris”: true/false,
open_doors_windows”: true/false,
site_accessible”: true/false,
exterior_mud”: true/false,
emergency_markings”: true/false,
major_repairs”: true/false,
vehicle_presence”: true/false
@@
{”house_destruction”: false,
structural_damage”: true,
exterior_debris”: true,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true}
A-A Prompts for decision stage (few-shot)
@@
house_destruction”: bool,
structural_damage”: bool,
exterior_debris”: bool,
open_doors_windows”: bool,
site_accessible”: bool,
exterior_mud”: bool,
emergency_markings”: bool,
major_repairs”: bool,
vehicle_presence”: bool
@@
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true
@@
house_destruction”: true,
structural_damage”: true,
exterior_debris”: true,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
@@
house_destruction”: false,
structural_damage”: true,
exterior_debris”: true,
open_doors_windows”: true,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true
@@
house_destruction”: false,
structural_damage”: true,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: true,
vehicle_presence”: false
@@
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: true,
major_repairs”: false,
vehicle_presence”: false
@@
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
@@
house_destruction”: true,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
@@
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: false,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
@@
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: true,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true
@@
{…JSON from vision model…}
@@
Output only one token: Occupied or Not Occupied’.
TABLE VII: Attribute schema used in VLM extraction: definitions and typical observable cues.
Attribute in prompts Typical visual cues (examples)
emergency_markings Official post-disaster markings: spray-painted search/inspection tags (e.g., X marks), placards, colored notices on doors/windows; exclude ordinary numbers or graffiti.

### v2

@@ Acknowledgments
Appendix A Supplementary Information
A.1 Prompts for VLM-only strategy
@@
"house_destruction": true/false,
"structural_damage": true/false,
"exterior_debris": true/false,
"open_doors_windows": true/false,
"site_accessible": true/false,
"exterior_mud": true/false,
"emergency_markings": true/false,
"major_repairs": true/false,
"vehicle_presence": true/false
@@
{"house_destruction": false,
"structural_damage": true,
"exterior_debris": true,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": true}
A.2 Prompts for decision stage (few-shot)
@@
"house_destruction": bool,
"structural_damage": bool,
"exterior_debris": bool,
"open_doors_windows": bool,
"site_accessible": bool,
"exterior_mud": bool,
"emergency_markings": bool,
"major_repairs": bool,
"vehicle_presence": bool
@@
"house_destruction": false,
"structural_damage": false,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": true
@@
"house_destruction": true,
"structural_damage": true,
"exterior_debris": true,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": false
@@
"house_destruction": false,
"structural_damage": true,
"exterior_debris": true,
"open_doors_windows": true,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": true
@@
"house_destruction": false,
"structural_damage": true,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": true,
"vehicle_presence": false
@@
"house_destruction": false,
"structural_damage": false,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": true,
"major_repairs": false,
"vehicle_presence": false
@@
"house_destruction": false,
"structural_damage": false,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": false
@@
"house_destruction": true,
"structural_damage": false,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": false
@@
"house_destruction": false,
"structural_damage": false,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": false,
"exterior_mud": false,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": false
@@
"house_destruction": false,
"structural_damage": false,
"exterior_debris": false,
"open_doors_windows": false,
"site_accessible": true,
"exterior_mud": true,
"emergency_markings": false,
"major_repairs": false,
"vehicle_presence": true
@@
{...JSON from vision model...}